### v1

# Sharp bounds on the half-space two-point function for high-dimensional Bernoulli percolation

Abstract
We consider Bernoulli percolation on $\mathbb{Z}^{d}$ with $d>6$. We prove an up-to-constant estimate for the critical two-point function restricted to a half-space. This completes previous results of Chatterjee and Hanson (Commun. Pure Appl. Math., 2021), and Chatterjee, Hanson, and Sosoe (Commun. Math. Phys., 2023), and solves a question asked by Hutchcroft, Michta, and Slade (Ann. Probab., 2023).
1 Introduction
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph with vertex set $\mathcal{V}$ and edge set $\mathcal{E}$. If $x,y\in\mathcal{V}$, we write $x\sim y$ to say that $\{x,y\}\in\mathcal{E}$. We consider Bernoulli percolation on $\mathcal{G}$. Given $p\in[0,1]$, we construct a random subgraph of $\mathcal{G}$ by independently keeping (resp. deleting) each edge of $\mathcal{E}$ with probability $p$ (resp. 1p1-p). The associated measure is denoted by $\mathbb{P}_{p}$. We focus on the following examples of graphs $\mathcal{G}$: for $d\geq 2$,
  1. $(i)$
    Nearest-neighbour model: $\mathcal{V}=\mathbb{Z}^{d}$ and $\mathcal{E}=\{\{x,y\}:\|x-y\|_{1}=1\}$ where $\|\cdot\|_{1}$ is the $\ell^{1}$ norm on $\mathbb{R}^{d}$;
  2. $(ii)$
    Spread-out model with spread parameter $L\geq 1$: $\mathcal{V}=\mathbb{Z}^{d}$ and $\mathcal{E}=\{\{x,y\}:\|x-y\|_{1}\leq L\}$.
It is well-known (see for instance [Gri99]) that the model undergoes a non-trivial phase transition as the parameter $p$ varies: one has $p_{c}\in(0,1)$ with
$p_{c}:=\inf\{p\in[0,1]:\mathbb{P}_{p}[0\xleftrightarrow{}\infty]>0\},$ (1.1)
where $\{0\xleftrightarrow{}\infty\}$ is the event that the origin lies in an infinite connected component.
In this paper, we study high-dimensional percolation, meaning that we work in dimensions $d>6$. This corresponds to the (conjectured) mean-field regime of the model. We refer to [Gri99, Sla06, Pan24, Hut25] and references therein for more information on the particular role played by the dimension $d=6$ in percolation theory. We investigate properties of the critical measure $\mathbb{P}=\mathbb{P}_{p_{c}}$. A fundamental quantity in its analysis is the so-called (restricted) two-point function, which is defined as follows: if $A\subset\mathbb{Z}^{d}$ and $x,y\in\mathbb{Z}^{d}$,
$\tau_{A}(x,y):=\mathbb{P}[x\xleftrightarrow{A\>}y],$ (1.2)
where $\{x\xleftrightarrow{A\>}y\}$ is the event that there exists an open path (i.e. a path made of edges that were kept) fully contained in $A$ which connects $x$ and $y$. When $A=\mathbb{Z}^{d}$, we drop it from the above notation.
The starting point in the study of high-dimensional Bernoulli percolation is the following estimate on the critical two-point function: for every $x,y\in\mathbb{Z}^{d}$,
$\tau(x,y)\asymp\frac{1}{1+|x-y|^{d-2}},$ (*)
where $\asymp$ means that the ratio of the two quantities is bounded away from 0 and infinity by two constants which only depend on $d$ (and potentially the spread parameter $L$), and where $|\cdot|$ denotes the $\ell^{\infty}$ norm on $\mathbb{R}^{d}$. The lace expansion approach developed by Brydges and Spencer [BS85] (see [Sla06] for a review) has been successfully implemented to derive a more precise version of (*1) for nearest-neighbour percolation in dimensions $d>10$ [HS90, Har08, FvdH17], and sufficiently spread-out percolation (i.e. $L\gg 1$) in dimensions $d>6$ [HHS03]. An alternative proof of (*1) in the latter setting has recently been obtained in [DCP25a].
We are interested in the behaviour of the critical two-point function restricted to the half-space $\mathbb{H}:=\{x=(x_{1},\dots,x_{d})\in\mathbb{Z}^{d}:x_{1}\geq 0\}$. In this setting, the main difficulty comes from the lack of full translation invariance. Nevertheless, several partial results have been obtained. The first set of results goes back to [CH20].
Proposition 1.1 ( [CH20]).
Let $d>6$ and assume that (*1) holds. Then, for every $K\geq 1$, there exist $c,C>0$ such that the following holds:
  1. $(a)$
    For every $x,y\in\mathbb{H}$ which satisfy $|x-y|\leq K\min(x_{1},y_{1})$,
    $\frac{c}{1+|x-y|^{d-2}}\leq\tau_{\mathbb{H}}(x,y)\leq\frac{C}{1+|x-y|^{d-2}}.$ (1.3)
  2. $(b)$
    For every $x,y\in\mathbb{H}$ which satisfy $x_{1}=0$ and $|x-y|\leq Ky_{1}$,
    $\frac{c}{1+|x-y|^{d-1}}\leq\tau_{\mathbb{H}}(x,y)\leq\frac{C}{1+|x-y|^{d-1}}.$ (1.4)
  3. $(c)$
    For every $x,y\in\mathbb{H}$ which satisfy $x_{1}=y_{1}=0$,
    $\frac{c}{1+|x-y|^{d}}\leq\tau_{\mathbb{H}}(x,y)\leq\frac{C}{1+|x-y|^{d}}.$ (1.5)
This result identifies three different regimes of decay for $\tau_{\mathbb{H}}(x,y)$. However, Proposition 1.1 does not give any information on how the two-point function interpolates between these regimes. Partial steps in this direction where taken in the subsequent work [CHS23]. Below, we let $\mathbf{e}_{1}=(1,0,\ldots,0)$.
Proposition 1.2 ( [CHS23]).
Let $d>6$ and assume that (*1) holds. Then, there exist $c,C>0$ such that the following holds:
  1. $(a)$
    For every $m\geq 1$, and every $x\in\mathbb{H}$,
    $\tau_{\mathbb{H}}(x,m\mathbf{e}_{1})\leq C\frac{1+m}{1+|x-y|^{d-1}}.$ (1.6)
  2. $(b)$
    For every $m\geq 1$, and every $x\in\mathbb{H}$, if $x_{1}\geq\tfrac{1}{2}|x|$ and $|x|\geq 4m$, then
    $\tau_{\mathbb{H}}(x,m\mathbf{e}_{1})\geq c\frac{1+m}{1+|x-y|^{d-1}}.$ (1.7)
Remark 1.3.
Some of the estimates stated in Propositions 1.1 and 1.2 have recently been derived in the context of spread-out percolation [DCP25a] (and also in the context of the weakly-self avoiding walk model [DCP25b]) using very different methods.
The estimates of Proposition 1.2 are inefficient in the situation were both $x$ and $y$ lie near the boundary of $\mathbb{H}$. By analogy with Green function estimates (see e.g [LL10]), Hutchcroft, Michta, and Slade [HMS23, Remark 3.4] conjectured a behaviour for $\tau_{\mathbb{H}}(x,y)$ in the regime where $\max(x_{1},y_{1})\leq|x-y|$. Our main result is a proof of their conjecture. It provides a sharp (up-to-constant) estimate on $\tau_{\mathbb{H}}(x,y)$ for every $x,y\in\mathbb{H}$. We will need the following notation: if $x,y\in\mathbb{H}$, we let $r_{x,y}:=\min(x_{1},|x-y|)$.
Theorem 1.4.
Let $d>6$ and assume that (*1) holds. Then, there exist $c,C>0$ such that, for every $x,y\in\mathbb{H}$,
$c\frac{(1+r_{x,y})\cdot(1+r_{y,x})}{1+|x-y|^{d}}\leq\tau_{\mathbb{H}}(x,y)\leq C\frac{(1+r_{x,y})\cdot(1+r_{y,x})}{1+|x-y|^{d}}.$ (1.8)
Remark 1.5.
It is interesting to compare our result with Proposition 1.2. The latter result can be rephrased as follows: if $d>6$ and (*1) holds, then there exist $c,C>0$ such that, for every $x,y\in\mathbb{H}$,
$\tau_{\mathbb{H}}(x,y)\leq C\frac{1+\min(r_{x,y},r_{y,x})}{1+|x-y|^{d-1}},$ (1.9)
and, assuming (for instance) that $x_{1}\geq\frac{1}{2}|x|$ and $|x|\geq 4|y|$,
$\tau_{\mathbb{H}}(x,y)\geq c\frac{1+y_{1}}{1+|x-y|^{d-1}}.$ (1.10)
Therefore, Theorem 1.4 corresponds to Proposition 1.2 in the regime where $r_{x,y}\asymp|x-y|$.
As an immediate corollary of Theorem 1.4, we obtain an alternative (short and easy) proof of [HMS23, Proposition 3.1] (which motivated [HMS23, Remark 3.4]). For every $n\in\mathbb{Z}$, let $\mathbb{H}_{n}:=\mathbb{H}-n\mathbf{e}_{1}$.
Corollary 1.6.
Let $d>6$ and assume that (*1) holds. Then, there exists $C>0$ such that, for every $n\geq 0$,
$\varphi_{p_{c}}(\mathbb{H}_{n}):=p_{c}\sum_{\begin{subarray}{c}x\in\mathbb{H}_{n}\\ y\notin\mathbb{H}_{n}\\ x\sim y\end{subarray}}\tau_{\mathbb{H}_{n}}(0,x)\leq C.$ (1.11)
Remark 1.7.
$(i$) This result was also derived in [DCP25a] in the context of spread-out percolation, and in [DCP25b] in the context of the weakly self-avoiding walk model.
$(ii)$ Corollary 1.6 implies the uniform boundedness of the expected number of critical pioneers of half-spaces.
Proof of Corollary 1.6.
By translation invariance, one has
$\varphi_{p_{c}}(\mathbb{H}_{n})=p_{c}\sum_{\begin{subarray}{c}x\in\mathbb{H}_{-n}\\ y\notin\mathbb{H}_{-n}\\ x\sim y\end{subarray}}\tau_{\mathbb{H}}(0,x).$ (1.12)
By Theorem 1.4, there exists $C_{1}>0$ such that, for every $x\in\mathbb{H}$
$\tau_{\mathbb{H}}(0,x)\leq C_{1}\frac{1+\min(x_{1},|x|)}{1+|x|^{d}}.$ (1.13)
Plugging (1.13) in (1.12) concludes the proof. ∎
Finally, let us mention that the upper bound in Theorem 1.4 is also useful in the recent [ASS25].
Notations.
We let $\|\cdot\|$ (resp. $|\cdot|$) denote the standard Euclidean norm (resp. the $\ell^{\infty}$ norm) on $\mathbb{R}^{d}$. If $f,g>0$, we write $f\lesssim g$ (or $g\gtrsim f$) if there exists $C>0$, which only depends on $d$ (and potentially the spread parameter $L$) such that $f\leq Cg$. If $f\lesssim g$ and $g\lesssim f$, we write $f\asymp g$.
Given $A,B,C\subset\mathbb{Z}^{d}$, we write $\{A\xleftrightarrow{C\>}B\}$ for the event that there exists an open path in $C$ connecting $A$ and $B$, and we omit the superscript $C$, when $C=\mathbb{Z}^{d}$.
We now introduce various “half-space notations”. Observe that these notations are slightly different from the standard ones. Given a subset $A\subset\mathbb{Z}^{d}$, we define the inner boundary of $A$ in $\mathbb{H}$,
$\partial A:=\{z\in A:\exists z^{\prime}\sim z\text{ with }z^{\prime}\in\mathbb{H}\cap A^{c}\},$ (1.14)
where we recall that $z^{\prime}\sim z$ means that $z$ and $z^{\prime}$ are neighbors in the graph $\mathcal{G}$ under consideration. For $z\in\mathbb{H}$, and $r\geq 0$, we denote the box of radius $r$ centered at $z$ in $\mathbb{H}$ as
$B_{r}(z)=\{y\in\mathbb{H}:|y-z|\leq r\},$ (1.15)
and just write $B_{r}$ when $z$ is the origin.
The van den Berg–Kesten inequality.
If $E$ and $F$ are two percolation events, we write $E\circ F$ for the event of disjoint occurrence of $E$ and $F$, that is, the event that there exist two disjoint sets $\mathcal{I}$ and $\mathcal{J}$ of edges such that the configuration restricted to $\mathcal{I}$ (resp. $\mathcal{J}$) is sufficient to decide that $E$ (resp. $F$) occurs. The van den Berg–Kesten (BK) inequality (see [Gri99, Section 2.3]) states that for two increasing events (i.e. events that are stable under the action of opening edges) $E$ and $F$, one has
$\mathbb{P}[E\circ F]\leq\mathbb{P}[E]\mathbb{P}[F].$ (BK)
2 Proof of Theorem 1.4
In the rest of the paper, we work either in the nearest-neighbour or in the spread-out setting (with spread parameter $L\geq 1$). Additionally, we assume that $d>6$ and that (*1) holds. The proof of Theorem 1.4 is based on Propositions 1.1 and 1.2 and on two new ingredients: Propositions 2.1 and 2.2. We state these results here and prove them in later sections.
We observe the following consequence of the BK inequality: for every $x,y\in\mathbb{H}$, letting $n=\lfloor|x-y|/3\rfloor$ and assuming that $n\geq 1$ (resp. $n\geq L$ in the spread-out case), one has
$\tau_{\mathbb{H}}(x,y)\leq\sum_{u\in\partial B_{n}(x)}\sum_{v\in\partial B_{n}(y)}\tau_{B_{n}(x)}(x,u)\cdot\tau_{\mathbb{H}}(u,v)\cdot\tau_{B_{n}(y)}(v,y).$ (2.1)
Indeed, exploring an open self-avoiding path $\gamma$ from $x$ to $y$, and decomposing it according to the last vertex $u$ visited by $\gamma$ before exiting $B_{n}(x)$, and the first vertex $v$ such that the restriction of $\gamma$ to the portion between $v$ and $y$ lies in $B_{n}(y)$ gives
$\{x\xleftrightarrow{\mathbb{H}\>}y\}\subset\bigcup_{u\in\partial B_{n}(x)}\bigcup_{v\in\partial B_{n}(y)}\{x\xleftrightarrow{B_{n}(x)\>}u\}\circ\{u\xleftrightarrow{\mathbb{H}\>}v\}\circ\{v\xleftrightarrow{B_{n}(y)\>}y\}.$ (2.2)
See Figure 1 for an illustration. Using a union bound and (BK) gives (2.1).
Figure 1: An illustration of the decomposition used to obtain (2.1) (in the nearest-neighbour case). The black bold path represents an open self-avoiding path from $x$ to $y$. Depending on the values of $x_{1}$ and $y_{1}$, the boxes $B_{n}(x)$ and $B_{n}(y)$ may “touch” the boundary of $\mathbb{H}$. The reversed inequality of Proposition 2.1 decomposes paths from $x$ to $y$ similarly, except that there is an additional restriction to vertices $u$ and $v$ satisfying $u_{1},v_{1}\geq\varepsilon n$.
Our first result provides a reversed inequality (see also Remark 2.4 below) up to some small multiplicative constant. Given $\varepsilon\in(0,1)$, $r\geq 0$ and $x\in\mathbb{H}$, we write
$\partial B_{r}^{\varepsilon}(x)=\{u\in\partial B_{r}(x):u_{1}\geq\varepsilon r\}.$ (2.3)
Proposition 2.1.
For every $\varepsilon\in(0,1/2)$, there exist $c,n_{0}>0$, such that for every $x,y\in\mathbb{H}$, letting $n=\lfloor|x-y|/3\rfloor$ and assuming that $n\geq n_{0}$, one has
$\tau_{\mathbb{H}}(x,y)\geq c\sum_{u\in\partial B_{n}^{\varepsilon}(x)}\sum_{v\in\partial B_{n}^{\varepsilon}(y)}\tau_{B_{n}(x)}(x,u)\cdot\tau_{\mathbb{H}}(u,v)\cdot\tau_{B_{n}(y)}(v,y).$ (2.4)
The second important new ingredient is the following estimate.
Proposition 2.2.
For every $n\geq 1$ and every $x\in\mathbb{H}$,
$\sum_{u\in\partial B_{n}(x)}\tau_{B_{n}(x)}(x,u)\lesssim\frac{1+\min(x_{1},n)}{n}.$ (2.5)
Furthermore, there exists $\varepsilon\in(0,1/2)$, such that, for every $n\geq 1$ and every $x\in\mathbb{H}$,
$\sum_{u\in\partial B_{n}^{\varepsilon}(x)}\tau_{B_{n}(x)}(x,u)\gtrsim\frac{1+\min(x_{1},n)}{n}.$ (2.6)
Remark 2.3.
$(1)$ The case $x_{1}=0$ of (2.5) was derived in [CH20, Lemma 26]. It is also known (see [HS14, Theorem 1.5]) that for every $n\geq 1$ and every $x\in\mathbb{H}$,
$\sum_{u\in\partial B_{n}(x)}\tau_{B_{n}(x)}(x,u)\lesssim 1.$ (2.7)
$(2)$ A weak version of (2.5) was derived in [DCP25a] in the context of (sufficiently) spread-out percolation. There, the authors obtained (see [DCP25a, Lemma 3.5]) the existence of $c_{0}=c_{0}(d)>0$ such that for every $n\geq 1$ and every $x\in\mathbb{H}$,
$\sum_{u\in\partial B_{n}(x)}\tau_{B_{n}(x)}(x,u)\lesssim\Big(\frac{1+\min(x_{1},n)}{n}\Big)^{c_{0}}.$ (2.8)
We postpone the proofs of these two propositions and give a short proof of our main result.
Proof of Theorem 1.4.
We begin with the upper bound. By (*1), one has, for every $u,v\in\mathbb{H}$,
$\tau_{\mathbb{H}}(u,v)\leq\tau(u,v)\lesssim\frac{1}{1+|u-v|^{d-2}}.$ (2.9)
Together with (2.1) and (2.5), this yields the existence of $C_{1}>0$ such that, for every $x,y\in\mathbb{H}$ with $n=\lfloor|x-y|/3\rfloor\geq 1$ (resp. $n\geq L$ in the spread-out case),
$\begin{split}\tau_{\mathbb{H}}(x,y)&\lesssim\frac{1}{n^{d-2}}\Big(\sum_{u\in\partial B_{n}(x)}\tau_{B_{n}(x)}(x,u)\Big)\cdot\Big(\sum_{v\in\partial B_{n}(y)}\tau_{B_{n}(y)}(y,v)\Big)\\ &\lesssim\frac{(1+r_{x,y})\cdot(1+r_{y,x})}{n^{d}}\leq C_{1}\frac{(1+r_{x,y})\cdot(1+r_{y,x})}{1+|x-y|^{d}},\end{split}$ (2.10)
where we used that $\min(x_{1},n)\lesssim r_{x,y}$ and $\min(y_{1},n)\lesssim r_{y,x}$. In the situation where $n=0$ (resp. $n<L$ in the spread-out case), a similar bound holds trivially (to the cost of potentially increasing $C_{1}$).
The lower bound follows similarly by combining Proposition 2.1 together with (1.4) and (2.6). Indeed, let $\varepsilon\in(0,1/2)$ be given by Proposition 2.2 and $n_{0}=n_{0}(\varepsilon)$ be given by Proposition 2.1, and observe that for every $x,y\in\mathbb{H}$ with $n=\lfloor|x-y|/3\rfloor\geq n_{0}$,
[…]
where we used that $\min(x_{1},n)\gtrsim r_{x,y}$ and $\min(y_{1},n)\gtrsim r_{y,x}$. Observe that if $u\in\partial B_{n}^{\varepsilon}(x)$ and $v\in\partial B_{n}^{\varepsilon}(y)$, then $\tfrac{n}{3}\leq|u-v|\leq\frac{10}{\varepsilon}\min(u_{1},v_{1})$. Using (1.3) with $K=10/\varepsilon$ gives $c_{1}=c_{1}(\varepsilon)$ such that,
$\min_{\begin{subarray}{c}u\in\partial B_{n}^{\varepsilon}(x)\\ v\in\partial B_{n}^{\varepsilon}(y)\end{subarray}}\tau_{\mathbb{H}}(u,v)\geq\frac{c_{1}}{n^{d-2}}.$ (2.12)
Plugging (2.12) in (2.11) gives $c_{2}>0$ such that, for every $x,y\in\mathbb{H}$ such that $n\geq n_{0}$,
$\tau_{\mathbb{H}}(x,y)\gtrsim\frac{(1+r_{x,y})\cdot(1+r_{y,x})}{n^{d}}\geq c_{2}\frac{(1+r_{x,y})\cdot(1+r_{y,x})}{1+|x-y|^{d}}.$ (2.13)
Again, the bound in the case $n\leq n_{0}$ follows straightforwardly (to the cost of potentially decreasing $c_{2}$). This concludes the proof. ∎
Remark 2.4.
Retrospectively, by combining (*1), Theorem 1.4, and (2.5), we can deduce that there exists some constant $c>0$ such that, for every $x,y\in\mathbb{H}$ with $n=\lfloor|x-y|/3\rfloor\geq 1$,
$\tau_{\mathbb{H}}(x,y)\geq c\sum_{u\in\partial B_{n}(x)}\sum_{v\in\partial B_{n}(y)}\tau_{B_{n}(x)}(x,u)\cdot\tau_{\mathbb{H}}(u,v)\cdot\tau_{B_{n}(y)}(v,y),$ (2.14)
which strengthens the result of Proposition 2.1.
3 Proof of Proposition 2.2
We now turn to the proof of Proposition 2.2.
Proof of (2.5).
Let $n\geq 1$. The case $x_{1}=0$ (resp. $x_{1}\leq L-1$ in the spread-out case) was derived [^2] For full disclosure, [CH20] only treats the case $x_{1}=0$. However, it is easy to extend their result to the case $x_{1}\leq L-1$ by using the Fortuin–Kasteleyn–Ginibre (FKG) inequality (see [Gri99, Chapter 2.2]). in [CH20, Lemma 26]. We now consider a general point $x\in\mathbb{H}$. Decomposing an open self-avoiding path $\gamma$ from $x$ to $u$ according to the earliest point $v\in\partial B_{n/4}(u)$ (along $\gamma$) such that the portion of $\gamma$ between $v$ and $u$ lies in $B_{n/4}(u)$ and using (BK), we obtain for any $u\in\partial B_{n}(x)$,
$\mathbb{P}[x\xleftrightarrow{B_{n}(x)\>}u]\leq\sum_{v\in\partial B_{n/4}(u)\cap B_{n}(x)}\mathbb{P}[x\xleftrightarrow{B_{n}(x)\>}v]\cdot\mathbb{P}[v\xleftrightarrow{B_{n/4}(u)\cap B_{n}(x)\>}u].$ (3.1)
Using (1.6) and (2.5) for $x_{1}=0$ (resp. $x_{1}\leq L-1$), we deduce that, for every $x\in\mathbb{H}$ and $u\in\partial B_{n}(x)$,
$\mathbb{P}[x\xleftrightarrow{B_{n}(x)\>}u]\lesssim\frac{1+\min(x_{1},n)}{n^{d-1}}\cdot\Big(\sum_{v\in\partial B_{n/4}(u)\cap B_{n}(x)}\mathbb{P}[v\xleftrightarrow{B_{n/4}(u)\cap B_{n}(x)\>}u]\Big)\lesssim\frac{1+\min(x_{1},n)}{n^{d}}.$ (3.2)
Summing over $u\in\partial B_{n}(x)$ concludes the proof. ∎
Proof of (2.6).
Let $\varepsilon\in(0,1/2)$ to be fixed. Let $n\geq 1$ and $x\in\mathbb{H}$. Without loss of generality, we may assume that $x=(x_{1},0,\ldots,0)$. We first assume that $n\geq\tfrac{1}{2}x_{1}$. Recall that $\mathbf{e}_{1}=(1,0,\ldots,0)$. On the one hand, by [^3] Let us give more details. If $\tfrac{1}{2}x_{1}\leq n\leq 2x_{1}$, then (1.3) (for a proper choice of $K$) gives (3.3). If $n\geq 2x_{1}$, we may apply (1.7) (since $x_{1}+2n\geq 4x_{1}$) to get (3.3). (1.3) and (1.7) , one has
$\tau_{\mathbb{H}}(x,x+2n\mathbf{e}_{1})\gtrsim\frac{1+\min(x_{1},n)}{n^{d-1}}.$ (3.3)
On the other hand, decomposing an open self-avoiding path from $x$ to $x+2n\mathbf{e}_{1}$ according to the first point in $\partial B_{n}(x)$ it visits and using (BK) gives
$\tau_{\mathbb{H}}(x,x+2n\mathbf{e}_{1})\lesssim\max_{u\in\partial B_{n}^{\varepsilon}(x)}\tau_{\mathbb{H}}(u,x+2n\mathbf{e}_{1})\cdot\Big(\sum_{u\in\partial B_{n}^{\varepsilon}(x)}\tau_{B_{n}(x)}(x,u)\Big)\\ +\max_{u\in\partial B_{n}(x)\setminus\partial B_{n}^{\varepsilon}(x)}\tau_{\mathbb{H}}(u,x+2n\mathbf{e}_{1})\cdot\Big(\sum_{u\in\partial B_{n}(x)\setminus\partial B_{n}^{\varepsilon}(x)}\tau_{B_{n}(x)}(x,u)\Big).$ (3.4)
Using (*1) and (1.6) give that, for $n$ large enough (in terms of $\varepsilon$),
$\max_{u\in\partial B_{n}^{\varepsilon}(x)}\tau_{\mathbb{H}}(u,x+2n\mathbf{e}_{1})\lesssim\frac{1}{n^{d-2}},\qquad\max_{u\in\partial B_{n}(x)\setminus\partial B_{n}^{\varepsilon}(x)}\tau_{\mathbb{H}}(u,x+2n\mathbf{e}_{1})\lesssim\frac{\varepsilon n}{n^{d-1}}.$ (3.5)
Combining (3.3), (3.4), and (3.5) gives, for $n$ large enough,
$\frac{1+\min(x_{1},n)}{n^{d-1}}\lesssim\frac{1}{n^{d-2}}\cdot\Big(\sum_{u\in\partial B_{n}^{\varepsilon}(x)}\tau_{B_{n}(x)}(x,u)\Big)+\varepsilon\cdot\frac{1+\min(x_{1},n)}{n^{d-1}},$ (3.6)
where we used (2.5) to get $\sum_{u\in\partial B_{n}(x)\setminus\partial B_{n}^{\varepsilon}(x)}\tau_{B_{n}(x)}(x,u)\lesssim\frac{1+\min(x_{1},n)}{n}$. Choosing $\varepsilon\in(0,1/2)$ small enough concludes the proof in the case $n\geq\tfrac{1}{2}x_{1}$ and $n$ large enough (in terms of $\varepsilon$). The remaining values of $n$ can be handled by adjusting the value of the constant in (2.6).
It remains to treat the case $n<\tfrac{1}{2}x_{1}$. Observe that for this choice, one has $B_{n}(x)=\{u\in\mathbb{Z}^{d}:|u-x|\leq n\}$ (that is, the $\mathbb{Z}^{d}$-box of radius $n$ around $x$ is fully included in $\mathbb{H}$). Thus, we may use [DCT16] to conclude that, for any $\varepsilon\in(0,1/2)$,
$\sum_{u\in\partial B_{n}^{\varepsilon}(x)}\tau_{B_{n}(x)}(x,u)\geq\frac{1}{2}\sum_{u\in\partial B_{n}(x)}\tau_{B_{n}(x)}(x,u)\gtrsim 1.$ (3.7)
This concludes the proof. ∎
4 Proof of Proposition 2.1
The proof of Proposition 2.1 is technically more involved. We will need a number of preliminary results and notations. In Section 4.1, we present a general lower bound for the probability that two sets are connected, which we express in terms of their (d4)(d-4)-capacity. In Section 4.2, we recall all the necessary definitions to introduce the theory of regularity of [KN11], taking here the viewpoint of [ASS25]. Finally, we give the proof of Proposition 2.1 in Section 4.3. Recall that we have assumed that $d>6$ and that (*1) holds.
4.1 A general lower bound for the connection probability of two finite sets
We present here a general lower bound for the probability that two finite subsets of $\mathbb{H}$ are connected by an open path in terms of the product of their (d4)(d-4)-capacity. A non-restricted version of this result has already appeared in [ASS25]. We adapt their argument to the setting of restricted percolation on $\mathbb{H}$. Since the proof in [ASS25] is only sketched, we provide here a more detailed argument for the reader’s convenience.
Recall that $\|\cdot\|$ denotes the Euclidean norm on $\mathbb{R}^{d}$. Given a non-empty finite subset $A\subset\mathbb{Z}^{d}$, we define its (d4)(d-4)-capacity as
$\textrm{Cap}_{d-4}(A):=\Big(\inf\Big\{\sum_{a,b\in A}\mu(a)\mu(b)(1+\|a-b\|)^{4-d}:\mu\textrm{ probability measure on }A\Big\}\Big)^{-1}.$ (4.1)
Given two finite sets $A,B\subset\mathbb{Z}^{d}$, we let ${\textup{d}}(A,B):=\min_{a\in A,b\in B}\|a-b\|$, and $\textrm{diam}(A):=\max_{a,a^{\prime}\in A}\|a-a^{\prime}\|$.
Lemma 4.1.
For every $c_{1}>0$, there exists $c>0$ such that the following holds. For every finite $A,B\subseteq\mathbb{H}$ such that $\textup{d}(A,B)\geq c_{1}\cdot\max(\textup{diam}(A),\textup{diam}(B))$, one has
$\mathbb{P}[A\xleftrightarrow{\mathbb{H}\>}B]\geq c\cdot\frac{\big(\min_{a\in A,b\in B}\tau_{\mathbb{H}}(a,b)\big)^{2}}{d(A,B)^{2-d}}\cdot\textup{Cap}_{d-4}(A)\cdot\textup{Cap}_{d-4}(B).$ (4.2)
Proof.
As for Lemma 8.1 in [ASS25], the proof is based on a second moment method. More precisely, given two probability measures $\mu$ and $\nu$ supported respectively on $A$ and $B$, consider the random variable
$X=\sum_{a\in A}\sum_{b\in B}\mu(a)\nu(b)\cdot\mathbf{1}\{a\xleftrightarrow{\mathbb{H}\>}b\}.$ (4.3)
Using Cauchy–Schwarz’s inequality, we get
$\mathbb{P}[A\xleftrightarrow{\mathbb{H}\>}B]\geq\mathbb{P}[X>0]\geq\frac{\mathbb{E}[X]^{2}}{\mathbb{E}[X^{2}]}.$ (4.4)
The first moment $\mathbb{E}[X]$ can be easily lower bounded: using that $\mu$ and $\nu$ are probability measures, we obtain that
$\mathbb{E}[X]\geq\min_{a\in A,b\in B}\tau_{\mathbb{H}}(a,b).$ (4.5)
We now upper bound $\mathbb{E}[X^{2}]$. Write
$\mathbb{E}[X^{2}]=\sum_{\begin{subarray}{c}a,a^{\prime}\in A\\ b,b^{\prime}\in B\end{subarray}}\mu(a)\mu(a^{\prime})\nu(b)\nu(b^{\prime})\mathbb{P}[a\xleftrightarrow{\mathbb{H}\>}b,a^{\prime}\xleftrightarrow{\mathbb{H}\>}b^{\prime}].$ (4.6)
Exploring an open self-avoiding path from $a$ to $b$, and then one from $a^{\prime}$ to $b^{\prime}$, and using (BK), we obtain that for every $a,a^{\prime}\in A$ and $b,b^{\prime}\in B$,
$\mathbb{P}[a\xleftrightarrow{\mathbb{H}\>}b,a^{\prime}\xleftrightarrow{\mathbb{H}\>}b^{\prime}]\leq\tau_{\mathbb{H}}(a,b)\tau_{\mathbb{H}}(a^{\prime},b^{\prime})+\sum_{w,w^{\prime}\in\mathbb{H}}\tau_{\mathbb{H}}(a,w)\tau_{\mathbb{H}}(a^{\prime},w)\tau_{\mathbb{H}}(w,w^{\prime})\tau_{\mathbb{H}}(w^{\prime},b)\tau_{\mathbb{H}}(w^{\prime},b^{\prime})\\ +\sum_{w,w^{\prime}\in\mathbb{H}}\tau_{\mathbb{H}}(a,w)\tau_{\mathbb{H}}(w,b^{\prime})\tau_{\mathbb{H}}(w,w^{\prime})\tau_{\mathbb{H}}(w^{\prime},a^{\prime})\tau_{\mathbb{H}}(w^{\prime},b).$ (4.7)
See Figure 2 for an illustration.
Figure 2: An illustration of the diagrams underlying the proof of (4.7). The black bold paths are open self-avoiding paths. If the event $\{a\xleftrightarrow{\mathbb{H}\>}b,a^{\prime}\xleftrightarrow{\mathbb{H}\>}b^{\prime}\}$ occurs, then one of the situations must occur (for some $w,w^{\prime}\in\mathbb{H}$). Each diagram corresponds to a term on the right-hand side of (4.7).
We begin with the analysis of the first sum on the right-hand side of (4.7). Letting $r=\textup{d}(A,B)/2$ and using (*1), we find that
$\begin{split}\Sigma(a,a^{\prime},b,b^{\prime})&:=\sum_{w,w^{\prime}\in\mathbb{H}}\tau_{\mathbb{H}}(a,w)\tau_{\mathbb{H}}(a^{\prime},w)\tau_{\mathbb{H}}(w,w^{\prime})\tau_{\mathbb{H}}(w^{\prime},b)\tau_{\mathbb{H}}(w^{\prime},b^{\prime})\\ &\lesssim\frac{1}{r^{d-2}}\Big(\sum_{w\in\mathbb{H}}\tau_{\mathbb{H}}(a,w)\tau_{\mathbb{H}}(a^{\prime},w)\Big)\cdot\Big(\sum_{w^{\prime}\in\mathbb{H}}\tau_{\mathbb{H}}(w^{\prime},b)\tau_{\mathbb{H}}(w^{\prime},b^{\prime})\Big)\\ &\qquad+\sum_{\begin{subarray}{c}w,w^{\prime}\in\mathbb{H}\\ \|w-w^{\prime}\|\leq r\end{subarray}}\tau_{\mathbb{H}}(a,w)\tau_{\mathbb{H}}(a^{\prime},w)\tau_{\mathbb{H}}(w,w^{\prime})\tau_{\mathbb{H}}(w^{\prime},b)\tau_{\mathbb{H}}(w^{\prime},b^{\prime}).\end{split}$ (4.8)
Using (*1) and a classical estimate (see for instance [DCP25a, Proposition B.1]),
$\sum_{w\in\mathbb{H}}\tau_{\mathbb{H}}(a,w)\tau_{\mathbb{H}}(a^{\prime},w)\leq\sum_{w\in\mathbb{Z}^{d}}\tau(a,w)\tau(w,a^{\prime})\lesssim(1+\|a-a^{\prime}\|)^{4-d},$ (4.9)
and likewise,
$\sum_{w^{\prime}\in\mathbb{H}}\tau_{\mathbb{H}}(w^{\prime},b)\tau_{\mathbb{H}}(w^{\prime},b^{\prime})\lesssim(1+\|b-b^{\prime}\|)^{4-d}.$ (4.10)
We now look at the second sum on the second line of (4.8). If $w$ is at distance at least $r/2$ from $A$, then,
$\sum_{\begin{subarray}{c}w^{\prime}\\ w:\textup{d}(w,A)\geq r/2,\>\|w-w^{\prime}\|\leq r\end{subarray}}\tau_{\mathbb{H}}(a,w)\tau_{\mathbb{H}}(a^{\prime},w)\tau_{\mathbb{H}}(w,w^{\prime})\tau_{\mathbb{H}}(w^{\prime},b)\tau_{\mathbb{H}}(w^{\prime},b^{\prime})\\ \lesssim\frac{1}{r^{2d-4}}\sum_{w^{\prime}}\tau_{\mathbb{H}}(w^{\prime},b)\tau_{\mathbb{H}}(w^{\prime},b^{\prime})\sum_{w:\|w-w^{\prime}\|\leq r}\tau(w,w^{\prime})\lesssim\frac{1}{r^{2d-6}}(1+\|b-b^{\prime}\|)^{4-d},$ (4.11)
where in the first inequality we used (*1), and in the second one we used (4.10) and (*1) one more time to get $\sum_{w:\|w-w^{\prime}\|\leq r}\tau(w,w^{\prime})\lesssim r^{2}$. Similarly,
$\sum_{\begin{subarray}{c}w\\ w^{\prime}:\textup{d}(w^{\prime},B)\geq r/2,\>\|w-w^{\prime}\|\leq r\end{subarray}}\tau_{\mathbb{H}}(a,w)\tau_{\mathbb{H}}(a^{\prime},w)\tau_{\mathbb{H}}(w,w^{\prime})\tau_{\mathbb{H}}(w^{\prime},b)\tau_{\mathbb{H}}(w^{\prime},b^{\prime})\lesssim\frac{1}{r^{2d-6}}(1+\|a-a^{\prime}\|)^{4-d}.$ (4.12)
By definition of $r$, it is impossible to have $\textup{d}(w,A)<r/2$, $\textup{d}(w^{\prime},B)<r/2$, and $\|w-w^{\prime}\|$ simultaneously. Hence,
$\sum_{\begin{subarray}{c}w,w^{\prime}\in\mathbb{H}\\ \|w-w^{\prime}\|\leq r\end{subarray}}\tau_{\mathbb{H}}(a,w)\tau_{\mathbb{H}}(a^{\prime},w)\tau_{\mathbb{H}}(w,w^{\prime})\tau_{\mathbb{H}}(w^{\prime},b)\tau_{\mathbb{H}}(w^{\prime},b^{\prime})\\ \lesssim\frac{1}{r^{2d-6}}\Big((1+\|a-a^{\prime}\|)^{4-d}+(1+\|b-b^{\prime}\|)^{4-d}\Big).$ (4.13)
Therefore, if we define for a probability measure $\rho$, its energy as
$\mathcal{E}_{d-4}(\rho):=\sum_{u,v\in\mathbb{Z}^{d}}\rho(u)\rho(v)(1+\|u-v\|)^{4-d},$ (4.14)
one obtains from the previously displayed equations that
$\begin{split}\sum_{\begin{subarray}{c}a,a^{\prime}\in A\\ b,b^{\prime}\in B\end{subarray}}\mu(a)\mu(a^{\prime})\nu(b)\nu(b^{\prime})\Sigma(a,a^{\prime},b,b^{\prime})&\lesssim\frac{1}{r^{d-2}}\mathcal{E}_{d-4}(\mu)\mathcal{E}_{d-4}(\nu)+\frac{1}{r^{2d-6}}\Big(\mathcal{E}_{d-4}(\mu)+\mathcal{E}_{d-4}(\nu)\Big)\\ &\lesssim\frac{1}{r^{d-2}}\cdot\mathcal{E}_{d-4}(\mu)\mathcal{E}_{d-4}(\nu),\end{split}$ (4.15)
where in the second inequality, we used that, since (by hypothesis) $r\gtrsim\max(\textrm{diam}(A),\textrm{diam}(B))$, one has $\mathcal{E}_{d-4}(\rho)\gtrsim\frac{1}{r^{d-4}}$, for any probability measure $\rho$ supported on $A$ or $B$.
Similar computations allow to treat the last sum in the right-hand side of (4.7). More precisely, using repeatedly (*1), we can see that it is upper bounded up to some multiplicative constant by
$\begin{split}&\frac{1}{r^{d-2}}\Big(\sum_{w\in\mathbb{H}}\tau_{\mathbb{H}}(a,w)\tau_{\mathbb{H}}(w,b^{\prime})\Big)\cdot\Big(\sum_{w^{\prime}\in\mathbb{H}}\tau_{\mathbb{H}}(w^{\prime},a^{\prime})\tau_{\mathbb{H}}(w^{\prime},b)\Big)\\ &\qquad\qquad+\sum_{\begin{subarray}{c}w,w^{\prime}\in\mathbb{H}\\ \|w-w^{\prime}\|\leq r\end{subarray}}\tau_{\mathbb{H}}(a,w)\tau_{\mathbb{H}}(w,b^{\prime})\tau_{\mathbb{H}}(w,w^{\prime})\tau_{\mathbb{H}}(w^{\prime},a^{\prime})\tau_{\mathbb{H}}(w^{\prime},b)\\ &\lesssim\frac{1}{r^{d-2}}\Big\{\frac{1}{r^{2(d-4)}}+\frac{1}{r^{d-2}}\sum_{\begin{subarray}{c}w,w^{\prime}\in\mathbb{H}\\ \|w-w^{\prime}\|\leq r\end{subarray}}\Big(\tau_{\mathbb{H}}(a,w)\tau_{\mathbb{H}}(w,w^{\prime})\tau_{\mathbb{H}}(w^{\prime},a^{\prime})+\tau_{\mathbb{H}}(w,b^{\prime})\tau_{\mathbb{H}}(w,w^{\prime})\tau_{\mathbb{H}}(w^{\prime},b)\Big)\Big\}\\ &\lesssim\frac{1}{r^{d-2}}\Big\{\mathcal{E}_{d-4}(\mu)\mathcal{E}_{d-4}(\nu)+\frac{1}{r^{d-2}}\Big(\frac{1}{1+\|a-a^{\prime}\|^{d-6}}+\frac{1}{1+\|b-b^{\prime}\|^{d-6}}\Big)\Big\}\\ &\lesssim\frac{1}{r^{d-2}}\Big\{\mathcal{E}_{d-4}(\mu)\mathcal{E}_{d-4}(\nu)+\frac{\mathcal{E}_{d-4}(\nu)}{1+\|a-a^{\prime}\|^{d-4}}+\frac{\mathcal{E}_{d-4}(\mu)}{1+\|b-b^{\prime}\|^{d-4}}\Big\},\end{split}$ (4.16)
where in the second inequality, we used (*1) to argue that
$\sum_{\begin{subarray}{c}w,w^{\prime}\in\mathbb{H}\\ \|w-w^{\prime}\|\leq r\end{subarray}}\tau_{\mathbb{H}}(a,w)\tau_{\mathbb{H}}(w,w^{\prime})\tau_{\mathbb{H}}(w^{\prime},a^{\prime})\leq\sum_{w,w^{\prime}\in\mathbb{Z}^{d}}\tau(a,w)\tau(w,w^{\prime})\tau(w^{\prime},a^{\prime})\lesssim\frac{1}{1+\|a-a^{\prime}\|^{d-6}}.$ (4.17)
Summing (4.16) over $a,a^{\prime}\in A$, and $b,b^{\prime}\in B$ against $\mu(a)\mu(a^{\prime})\nu(b)\nu(b^{\prime})$, we obtain the same upper bound as in (4.15), and conclude that
$\mathbb{E}[X^{2}]\lesssim\frac{1}{r^{d-2}}\cdot\mathcal{E}_{d-4}(\mu)\mathcal{E}_{d-4}(\nu).$ (4.18)
Optimising over the choices of $\mu$ and $\nu$ and combining the result with (4.4) and (4.5) concludes the proof. ∎
4.2 Regular points, line good points, extended cluster
We present here the basis of a technique first introduced in [KN11] to derive the one-arm exponent in high-dimensional critical percolation. It is based on a notion of regularity. Here, we will adapt to our setting the definition of regular points from [ASS25] which is stated in purely geometric terms.
We write $\mathcal{C}(x;A):=\{z\in A:x\stackrel{{\scriptstyle A}}{{\longleftrightarrow}}z\}$ for the cluster of a point $x$ restricted to a set $A$. Fix $n\geq 1$ and $x\in\mathbb{H}$, and for $z\in\partial B_{n}(x)$, and $s>0$, consider the event
$\mathcal{T}_{s}(z):=\{|\mathcal{C}(z;B_{n}(x))\cap B(z,s)|\leq s^{4}(\log s)^{7}\}\\ \cap\{|\mathcal{C}(z;B_{n}(x))\cap B(z,s)\cap\partial B_{n}(x)|\leq s^{2}(\log s)^{7}\}.$ (4.19)
Definition 4.2 ($K$-regular points).
Given $K>0$, we call $z\in\partial B_{n}(x)$ a $K$-regular point, if the events $\mathcal{T}_{s}(z)$ hold for all $s\geq K$.
Let $\varepsilon\in[0,1/2)$, We denote by $X_{n}^{\varepsilon,K-\textup{reg}}(x)$ the number of points on $\partial B_{n}^{\varepsilon}(x)$ (recall (2.3)), which are $K$-regular and connected to $x$ in $B_{n}(x)$. Also, denote by $X_{n}^{\varepsilon}(x)$ the number of points on $\partial B_{n}^{\varepsilon}(x)$ which are connected to $x$ in $B_{n}(x)$ (the so-called pioneers). It turns out that most of the pioneers are regular, and consequently one can show the following lemma.
Lemma 4.3.
There exist $K_{0}\geq 1$ and $n_{0}\geq 1$ such that the following holds. For every $K\geq K_{0}$, every $n\geq n_{0}$, every $\varepsilon\in[0,1/2)$, and every $x\in\mathbb{H}$,
$\mathbb{E}[X_{n}^{\varepsilon,K-\textup{reg}}(x)]\geq\frac{1}{2}\cdot\mathbb{E}[X_{n}^{\varepsilon}(x)].$ (4.20)
We defer the proof of this lemma to Section 4.4 and introduce now the notion of $K$-line good points. For this, one first needs to consider a maximal subset of the set $X_{n}^{\varepsilon,K-\textup{reg}}(x)$ of $K$-regular points, which has the property that all its points are at distance at least $2K$ one from each other. Denote by $\mathcal{X}_{n}^{\varepsilon,K-\textrm{reg}}(x)$ one such maximal subset chosen uniformly at random. Then, if $z\in\mathcal{X}_{n}^{\varepsilon,K-\textrm{reg}}(x)$, we consider a line segment of length $K$ emanating from $z$, outside $B_{n}(x)$ and orthogonal to its boundary (choose one arbitrarily if there are many). Call $z^{\prime}$ the endpoint of this line segment. We say that $z^{\prime}$ is a $K$-line good point if all the edges on the line segment between $z$ and $z^{\prime}$ are open. More generally, for any $z\in\mathcal{X}_{n}^{\varepsilon,K-\textrm{reg}}(x)$, we denote by $L_{z}$ the maximal open segment emanating from $z$ orthogonally to $B_{n}(x)$, of length at most $K$.
We next define the extended cluster of $x$ in $B_{n}(x)$, which we denote by $\mathcal{C}_{n}^{e}(x)$, as the cluster of $x$ in $B_{n}(x)$ together with all the line segments $L_{z}$ for $z\in\mathcal{X}_{n}^{\varepsilon,K-\textrm{reg}}(x)$.
We say that a set $A$ is $K$-admissible for the pair $(x,n)$, if $\mathbb{P}[\mathcal{C}_{n}^{e}(x)=A]>0$, and for such admissible set we denote by $\partial_{*}A$ its set of points which are at distance exactly $K$ from $B_{n}(x)$. Hence by definition $\partial_{*}\mathcal{C}_{n}^{e}(x)$ is the set of $K$-line good points.
One interest of the notion of regularity, which has been noticed and used extensively in [ASS25], is that in any dimension $d>6$, admissible sets have a (d4)(d-4)-capacity which is comparable to their cardinality. Indeed, the following lemma was observed in [ASS25, Claim 6.1].
Lemma 4.4.
There exists a constant $c>0$, such that for every $n,K\geq 1$, every $\varepsilon\in[0,1/2)$, every $x\in\mathbb{H}$, and every $K$-admissible set $A$ for the pair $(x,n)$, one has
$\textup{Cap}_{d-4}(\partial_{*}A)\geq c|\partial_{*}A|.$ (4.21)
Proof.
For the reader’s convenience, we include a short proof. By taking $\mu$ to be the uniform measure on $\partial_{*}A$ in the definition of the (d4)(d-4)-capacity, one gets
$\textrm{Cap}_{d-4}(\partial_{*}A)\geq\frac{|\partial_{*}A|^{2}}{\sum_{a,a^{\prime}\in\partial_{*}A}(1+\|a-a^{\prime}\|)^{4-d}}.$ (4.22)
Now, by definition of a $K$-regular point, one has for any fixed $a\in\partial_{*}A$,
$\sum_{a^{\prime}\in\partial_{*}A}(1+\|a-a^{\prime}\|)^{4-d}\lesssim\sum_{i\geq\log_{2}(K)}\frac{2^{2i}\cdot i^{7}}{2^{(d-4)i}}\lesssim 1,$ (4.23)
where the implicit constants do not depend on $x$, $n$, $K$, $\varepsilon$, and $A$. This concludes the proof. ∎
4.3 Conclusion
We now have all the necessary material to prove our desired result.
Proof of Proposition 2.1.
Let $\varepsilon\in(0,1/2)$. Let $x,y\in\mathbb{H}$ and set $n=\lfloor|x-y|/3\rfloor$. Fix $K_{0},n_{0}$ as in Lemma 4.3. Let $K\geq K_{0}$ to be chosen large enough and assume that $n\geq n_{0}$.
We first observe that
$\tau_{\mathbb{H}}(x,y)\geq\sum_{A,B}\mathbb{P}\big[\mathcal{C}_{n}^{e}(x)=A,\mathcal{C}_{n}^{e}(y)=B,\partial_{*}A\xleftrightarrow{\textrm{off }(A\cup B)\>}\partial_{*}B\big],$ (4.24)
where $\{\partial_{*}A\xleftrightarrow{\textrm{off }(A\cup B)\>}\partial_{*}B\}$ is the event that $\partial_{*}A$ is connected to $\partial_{*}B$ by an open path in $\mathbb{H}$ that avoids $A\cup B$, except at its end points. Note next that the two events $\{\mathcal{C}_{n}^{e}(x)=A,\mathcal{C}_{n}^{e}(y)=B\}$ and $\{\partial_{*}A\xleftrightarrow{\textrm{off }(A\cup B)\>}\partial_{*}B\}$ depend on different sets of edges, and are thus independent. Hence, we get
$\tau_{\mathbb{H}}(x,y)\geq\sum_{A,B}\mathbb{P}\big[\mathcal{C}_{n}^{e}(x)=A,\mathcal{C}_{n}^{e}(y)=B\big]\cdot\mathbb{P}\big[\partial_{*}A\xleftrightarrow{\textrm{off }(A\cup B)\>}\partial_{*}B\big].$ (4.25)
Given $A,B$ which are $K$-admissible respectively for $(x,n)$ and $(y,n)$, we define
$C=A\cup B\cup\big(\bigcup_{a\in\partial_{*}A}B_{K}(a)\big)\cup\big(\bigcup_{b\in\partial_{*}B}B_{K}(b)\big).$ (4.26)
Now, we fix an arbitrary ordering of the elements of $\partial_{*}A$ and $\partial_{*}B$. On the event $\{\partial_{*}A\xleftrightarrow{\textrm{off }(A\cup B)\>}\partial_{*}B\}$, we denote by $Y_{1}$ the first element $a\in\partial_{*}A$ for this ordering such that $\partial B_{K}(a)$ is connected to $\cup_{b\in\partial_{*}B}B_{K}(b)$ by an open path that avoids $C$, and let $Y_{2}$ be the first element $b\in\partial_{*}B$ such that $\partial B_{K}(Y_{1})$ is connected to $\partial B_{K}(b)$ by an open path that avoids $C$. Finally, we let
$H=(A\cup B)\setminus\big(\bigcup_{z\in\partial_{*}A\cup\partial_{*}B}L_{z}\big),$ (4.27)
where for any $a\in\partial_{*}A$ we let $L_{a}$ be the line segment of length $K$ between $a$ and $B_{n}(x)$, and similarly for $b\in\partial_{*}B$. Letting $c(K)>0$ be the probability that all edges are open in a box of size $K$, we find
[…]
We then write
$\mathbb{P}\big[\partial_{*}A\xleftrightarrow{\textrm{off }H\>}\partial_{*}B\big]=\mathbb{P}\big[\partial_{*}A\xleftrightarrow{}\partial_{*}B\big]-\mathbb{P}\big[\partial_{*}A\xleftrightarrow{\textrm{via }H\>}\partial_{*}B\big],$ (4.29)
where $\{\partial_{*}A\xleftrightarrow{\textrm{via }H\>}\partial_{*}B\}$ denotes the event that $\partial_{*}A$ and $\partial_{*}B$ are connected by an open path, and all open paths that connect them intersect $H$. We claim that for any constant $\delta>0$, one can find $K\geq K_{0}$ large enough, so that for all admissible sets $A$ and $B$,
$\mathbb{P}\big[\partial_{*}A\xleftrightarrow{\textrm{via }H\>}\partial_{*}B\big]\leq\frac{\delta}{n^{d-2}}\cdot|\partial_{*}A|\cdot|\partial_{*}B|.$ (4.30)
To see this, we note that by a union bound, it suffices to show that for $K$ sufficiently large, for every $a\in\partial_{*}A$ and $b\in\partial_{*}B$, one has
$\mathbb{P}\big[a\xleftrightarrow{\textrm{via }H\>}b\big]\leq\frac{\delta}{n^{d-2}}.$ (4.31)
Now, decomposing an open self-avoiding path connecting $a$ and $b$ through $H$ according to the first point in $H$ it visits, and using (BK) and (*1), we deduce that
$\begin{split}\mathbb{P}\big[a\xleftrightarrow{\textrm{via }H\>}b\big]&\leq\sum_{u\in H}\tau_{\mathbb{H}}(a,u)\tau_{\mathbb{H}}(u,b)\lesssim\frac{1}{n^{d-2}}\Big(\sum_{u\in H\cap A}\tau_{\mathbb{H}}(a,u)+\sum_{u\in H\cap B}\tau_{\mathbb{H}}(u,b)\Big)\\ &\lesssim\frac{1}{n^{d-2}}\sum_{i\geq\log_{2}(K)}\frac{2^{4i}\cdot i^{7}}{2^{(d-2)i}}\lesssim\frac{1}{n^{d-2}\sqrt{K}},\end{split}$ (4.32)
which gives (4.31)—and hence (4.30)—by choosing $K$ large enough.
On the other hand, by combining Lemmas 4.1 and 4.4, we obtain that for every $K$-admissible sets $A$ and $B$,
$\mathbb{P}\big[\partial_{*}A\longleftrightarrow\partial_{*}B\big]\gtrsim\Big(\min_{\begin{subarray}{c}a\in\partial_{*}A\\ b\in\partial_{*}B\end{subarray}}\tau_{\mathbb{H}}(a,b)\Big)^{2}\cdot n^{d-2}\cdot|\partial_{*}A|\cdot|\partial_{*}B|.$ (4.33)
If $a\in\partial_{*}A$ (resp. $b\in\partial_{*}B$), then $a_{1}\geq\varepsilon n$ (resp. $b_{1}\geq\varepsilon n$). As a result, for $(a,b)\in\partial_{*}A\times\partial_{*}B$, one has $n\lesssim|a-b|\lesssim\min(a_{1},b_{1})$ (where the implicit constants depend on $\varepsilon$). This is where we crucially require that $\varepsilon>0$. We can therefore use (1.3) to conclude that
$\Big(\min_{\begin{subarray}{c}a\in\partial_{*}A\\ b\in\partial_{*}B\end{subarray}}\tau_{\mathbb{H}}(a,b)\Big)^{2}\gtrsim(n^{2-d})^{2}.$ (4.34)
Altogether, this shows that, with the notation of Lemma 4.3,
$\begin{split}\tau_{\mathbb{H}}(x,y)&\gtrsim\frac{1}{n^{d-2}}\sum_{A,B}\mathbb{P}\big[\mathcal{C}_{n}^{e}(x)=A,\mathcal{C}_{n}^{e}(y)=B\big]\cdot|\partial_{*}A|\cdot|\partial_{*}B|\\ &=\frac{1}{n^{d-2}}\cdot\mathbb{E}\big[|\partial_{*}\mathcal{C}_{n}^{e}(x)|\big]\cdot\mathbb{E}\big[|\partial_{*}\mathcal{C}_{n}^{e}(y)|\big]\\ &\gtrsim\frac{1}{n^{d-2}}\cdot\mathbb{E}\big[X_{n}^{\varepsilon,K-\textrm{reg}}(x)\big]\cdot\mathbb{E}\big[X_{n}^{\varepsilon,K-\textrm{reg}}(y)\big]\\ &\gtrsim\frac{1}{n^{d-2}}\mathbb{E}[X_{n}^{\varepsilon}(x)]\cdot\mathbb{E}[X_{n}^{\varepsilon}(y)],\end{split}$ (4.35)
where we used Lemma 4.3 in the last inequality. Using (*1) in (4.35) yields the existence of $c=c(\varepsilon)>0$ such that, for every $x,y\in\mathbb{H}$ satisfying $\lfloor|x-y|/3\rfloor\geq n_{0}$,
$\tau_{\mathbb{H}}(x,y)\geq c\sum_{u\in\partial B_{n}^{\varepsilon}(x)}\sum_{v\in\partial B_{n}^{\varepsilon}(y)}\tau_{B_{n}(x)}(x,u)\cdot\tau_{\mathbb{H}}(u,v)\cdot\tau_{B_{n}(y)}(v,y).$ (4.36)
This concludes the proof. ∎
4.4 Proof of Lemma 4.3
The proof of Lemma 4.3 is very similar to the proofs of Theorem 4 in [KN11] and Proposition 5.7 in [ASS25]. First, one needs to introduce a local density condition. To be more precise, fix $n\geq 1$, $x\in\mathbb{H}$, and for $s>0$ and $z\in\partial B_{n}(x)$, consider the event
[…]
The interest of this event, when compared to $\mathcal{T}_{s}(z)$, is that it only depends on the configuration of the percolation inside the box $B(z,s^{d})$, and is thus a purely local event, while to determine whether $\mathcal{T}_{s}(z)$ holds or not, one needs a priori to know the configuration in the whole box $B_{n}(x)$. The drawback is that this event is a priori less likely than $\mathcal{T}_{s}(z)$, but as Lemma 4.6 below shows, it is still extremely likely, and furthermore, the following simple fact holds by construction (see Claim 4.1 in [KN11] or Claim 5.4 in [ASS25]).
Lemma 4.5.
One has for every $n\geq 1$, every $x\in\mathbb{H}$, every $z\in\partial B_{n}(x)$, and every $s>0$,
$\mathcal{T}_{s}^{\textup{loc}}(z)\subseteq\mathcal{T}_{s}(z).$ (4.38)
As already mentioned, another fact we will use, and which is proved in [ASS25, Claim 5.5], is the following.
Lemma 4.6.
There exists a constant $c>0$, such that for every $n\geq 1$, every $x\in\mathbb{H}$, every $z\in\partial B_{n}(x)$, and every $s>0$,
$\mathbb{P}[\mathcal{T}_{s}^{\textup{loc}}(z)]\geq 1-\exp\big(-c(\log s)^{4}\big).$ (4.39)
The last fact we shall need is Lemma 1.1 from [KN11], which we state here for the sake of completeness.
Lemma 4.7.
There exist positive constants $c$ and $C$, such that for every $u\in\mathbb{H}$, every $s>0$, and every $z\in\partial B_{s}(u)$,
$\tau_{B_{s}(u)}(u,v)\geq c\exp(-C(\log s)^{2}).$ (4.40)
Using the FKG inequality, it follows from Lemma 4.7, that the probability to connect two arbitrary points of $\partial B_{s}(u)$ (for some $u\in\mathbb{H}$) is at least $c^{2}\exp(-2C(\log s)^{2})$. We are now in a position to prove Lemma 4.3.
Proof of Lemma 4.3.
Fix $x\in\mathbb{H}$ and $\varepsilon\in(0,1)$. Let $n,K\geq 1$ to be fixed. Let us say that a point $z\in\partial B_{n}(x)$ is $s$-locally bad if the event $\mathcal{T}_{s}^{\textup{loc}}(z)$ does not hold, and let us denote by $X_{n}^{\varepsilon,s\textup{-loc-bad}}(x)$ the number of points on $\partial B_{n}^{\varepsilon}(x)$ which are $s$-locally bad and connected to $x$ in $B_{n}(x)$. Note that, due to Lemma 4.5, one has
$\mathbb{E}[X_{n}^{\varepsilon,K-\textrm{reg}}(x)]\geq\mathbb{E}[X_{n}^{\varepsilon}(x)]-\sum_{s\geq K}\mathbb{E}[X_{n}^{\varepsilon,s\textup{-loc-bad}}(x)].$ (4.41)
We now upper bound each term of the above sum separately. First of all, for $s$ such that $8s^{d}>n$, we simply use Lemma 4.6 and a union bound over all the points on $\partial B_{n}(x)$, to get (for some $c_{1}=c_{1}(d)>0$)
$\mathbb{E}[X_{n}^{\varepsilon,s\textup{-loc-bad}}(x)]\lesssim n^{d-1}\cdot\exp(-c_{1}(\log n)^{4}).$ (4.42)
Additionally, for $s$ such that $s^{2}$ is much larger than $n^{d-1}$ (or larger than the total number of points on $\partial B_{n}(x)$), the set of $s$-locally bad points is empty by definition so that $\mathbb{E}[X_{n}^{\varepsilon,s\textup{-loc-bad}}(x)]=0$. On the other hand one has by Lemma 4.7 (and the remark following it),
$\mathbb{E}[X_{n}^{\varepsilon}(x)]\gtrsim\exp(-C_{1}(\log n)^{2}),$ (4.43)
for some constant $C_{1}>0$. Combining the two previously displayed equations, one can choose $n$ large enough so that
$\sum_{s:8s^{d}>n}\mathbb{E}[X_{n}^{\varepsilon,s\textup{-loc-bad}}(x)]\leq\frac{1}{4}\cdot\mathbb{E}[X_{n}^{\varepsilon}(x)].$ (4.44)
We now fix $s\geq K$ such that $8s^{d}\leq n$ and consider the set $U=\{u\in\mathbb{Z}^{d}:u_{i}\in\{0,s^{d}\}\ \forall i=1,\dots,d\}$. For each $w\in U$, we define
$\mathcal{B}(w)=\{B(z,2s^{d}):z\in w+4s^{d}\cdot\mathbb{Z}^{d}\}.$ (4.45)
Denote by $Q(w)$ the union of all the boxes of this partition which intersect $\partial B_{n}^{\varepsilon}(x)$, and let $N(w)$ be the number of those which are connected to $x$ inside $B_{n}(x)\setminus Q(w)$. By Lemma 4.7 and the remark following it, one has $c_{2},C_{2}>0$ such that, for any $w\in U$,
$\mathbb{E}[X_{n}^{\varepsilon}(x)]\geq c_{2}\exp(-C_{2}(\log s)^{2})\cdot\mathbb{E}[N(w)],$ (4.46)
and hence also (since $|U|=2^{d}$),
$\mathbb{E}[X_{n}^{\varepsilon}(x)]\geq\frac{c_{2}}{2^{d}}\exp(-C_{2}(\log s)^{2})\cdot\sum_{w\in U}\mathbb{E}[N(w)].$ (4.47)
For a box $q\in Q(w)$, call the interior of $q$ the set of points in $q$ which are at distance at least $s^{d}$ from the points which are in $q^{c}\cap B_{n}(x)$. Observe that as $w$ varies in $U$, the union of all the interiors of the boxes $q\in Q(w)$ covers the whole boundary $\partial B_{n}^{\varepsilon}(x)$ (recall that we assume $8s^{d}\leq n$). Note also that for a point $z$ on $\partial B_{n}^{\varepsilon}(x)$ which is in the interior of a box $q\in Q(w)$, the event $\mathcal{T}_{s}^{\textup{loc}}(z)$ only depends on the configuration of edges inside $q$. Since there are at most order $s^{d^{2}}$ such points in each box $q\in Q(w)$, a union bound and Lemma 4.6 give that for some constants $c_{3},C_{3}>0$, and for any $s$ as above,
$\mathbb{E}[X_{n}^{\varepsilon,s\textup{-loc-bad}}(x)]\leq C_{3}s^{d^{2}}\exp(-c_{3}(\log s)^{4})\sum_{w\in U}\mathbb{E}[N(w)].$ (4.48)
Hence, using again (4.47), and taking $K$ large enough ensures that for all $n\geq 1$,
$\sum_{s:K\leq s\leq(n/8)^{1/d}}\mathbb{E}[X_{n}^{\varepsilon,s\textup{-loc-bad}}(x)]\leq\frac{1}{4}\cdot\mathbb{E}[X_{n}^{\varepsilon}(x)].$ (4.49)
Together with (4.41) and (4.44), this concludes the proof of the lemma. ∎
Acknowledgements.
RP thanks Hugo Duminil-Copin for stimulating discussions at an early stage of this project, and acknowledges the support of the Swiss National Science Foundation through a Postdoc.Mobility grant. BS acknowledges the support from the grant ANR-22-CE40-0012 (project LOCAL).